[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 25, 2016

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Wed May 25 11:39:05 UTC 2016


Update..To be more accurate. NCO permanently solved the networking issue
with the comp server 03.
Thanks
Chris

On Wed, May 25, 2016 at 7:29 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:

> Here is today's summary of NHC computer operations:
> --- mello
>
> ----------
>
> *TAFB:*
>
> TAFB noted an isolated problem sending the GFE grids to NAWIPS Monday.
> Networking issues with the comp server 03 caused the problem. NCO is
> working to ensure the problem does not occur again.
> *HSU:*
>
> TSB fixed the problem with the UKMET data flow. Consequently, it also
> fixed the GUNC model (The GUNC still lives !).
> *Other items:*
>
> The Himawari data is flowing again.
> *NHC RFCs:*
>
>
>


-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160525/255dc279/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list